OpenStudy (anonymous):

Your Firm offers to deliver 300 tables to a dealer, at $90 per table, and to reduce the price per table on the entire order by 25 cents for each additional table over 300. a) find the dollar total involved in the largest possible possible transaction between you and the dealer. b)Find the smallest possible dollar amount.

OpenStudy (anonymous):

i am not completely sure but i think it should be y = 300*90 -x*.25

OpenStudy (anonymous):

where x is the table count greater that 300

OpenStudy (anonymous):

since it is a linear equation the max money transfer is when x = 0, so 300*90

OpenStudy (anonymous):

and for the min is when y = 0; anyone else have an input on this?

OpenStudy (anonymous):

Surely the linear equation should be \[y = 90(300+x) - \frac{1}{4}x = 27000 + 89.75x\] since they still have to pay for each table over 300, but get an additional discount also.